Sheridan Music Company makes custom marimbas and xylophones. Since much of the work on these musical instruments is done by hand, the company uses direct labor hours as its manufacturing overhead application base. The company’s annual budgeted overhead costs for 72,600 direct labor hours totaled $580,800

. (a) Assume that during the year, the company incurred manufacturing overhead totaling $580,400 for 72,600 direct labor hours. By how much was manufacturing overhead under- or overapplied for the year? (Enter negative amount using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Overhead $ ??

(b) Assume that during the year, the company incurred manufacturing overhead totaling $580,400 for 75,350 direct labor hours. By how much was manufacturing overhead under- or overapplied for the year? (Enter negative amount using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Overhead $ ??

(c) Assume that during the year, the company incurred manufacturing overhead totaling $605,000 for 75,350 direct labor hours. By how much was manufacturing overhead under- or overapplied for the year? (Enter negative amount using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Overhead $??
Answer #1

Overhead rate = 580800/72600 = 8 per DLH

a) Overhead applied = 72600*8 = 580800

Actual overhead = 580400

Over applied overhead = 580800-580400 = 400

b) Overhead applied = 75350*8 = 602800

Actual overhead = 580400

Over applied overhead = 602800-580400 = 22400

c) Overhead applied = 602800

Actual overhead = 605000

Under applied overhead = 602800-605000 = -2200